22west:

Which expression shows the result of applying the distributive property to 3(2x−6) ? 5x−3 6x−6 2x−18 6x−18

dude:

Do you know what the distribute property is/says?

22west:

kind of

dude:

Here is an example Use distributive property on this 1(x+3)

22west:

ok

dude:



22west:

where did you get the 1

dude:

From my sample question

22west:

oh

22west:

3*x , 3*2, and 3*6

dude:

Ill focus on yours 

22west:

oh

22west:

A. 5x−3 B. 6x−6 C. 2x−18 D. 6x−18

22west:

do i add that now

dude:

My drawing is terrible but you multiply the outer part of the parenthesis into the inside 3(2x-6) gives \(3\cdot 2x~~and ~~3\cdot -6\)

22west:

ok so do i add that

dude:

Multiply first and then add XD

22west:

3*2=6 3*(-6)=-18

dude:

Yep, so we have \(6x-18\)

22west:

ok

dude:

That is your answer XD

22west:

i have some more

dude:

Sure, make a new post

22west:

ok